# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ 5.3 Experience Replay
Q-learning is facing the problem that creating trajectories and samples can be a time-consuming task, especially if the dynamics of the system are difficult to simulate or the time horizon is large. This makes it intractable to use generated samples only once and then throw them away when the parameter update is complete. We will use a technique called experience replay LinLong-Ji ; Mnih2013 that solves the problem of wasting generated samples. This uses a so-called replay memory or replay buffer $\pazocal{R}$ with size $\pazocal{N}_{R}\in\mathbb{N}$ to store samples $(x_{n},a_{n},x_{n+1})\in\pazocal{R}$ for $n=0,\ldots,N-1$ and replay (reuse) them as needed in batches $(x_{n}^{j},a_{n}^{j},x_{n+1}^{j}),j=1,\ldots,M$ to update the parameters. Sampling past experience, for example, from a uniform distribution on $\pazocal{R}$, also helps to overcome the exploration-exploitation dilemma. Here, the samples obtained in the early phase of the algorithm are used repeatedly, making the choice of the exploration rate less important for the performance of the algorithm. We summarize the Q-learning method with replay buffer in Algorithm 2.
@@
Step 1: Initialize $(\theta_{n}^{0})_{n=0}^{N-1}=(\theta_{n}^{0})_{n=0}^{N-1}$; set the maximum number of iterations $k^{\max}>0$, batch size $M$; choose a selection policy $u^{S}=(u_{n}^{S})_{n=0}^{N-1}$
Step 2: Set $n=0$; choose the initial states $x_{0}\in\pazocal{X}$
@@
     Choose $\alpha_{n}^{k}\in[0,1]$ and update parameters: $\theta_{n}^{k+1}=\theta^{k}_{n}-\alpha_{n}^{k}\frac{1}{M}\sum_{j=1}^{M}\delta_{n}^{j}\nabla_{\theta_{n}}\overline{Q}_{n}(x_{n}^{j},a_{n}^{j},\theta_{n}^{k})$
@@
Step 3: Set $k=k+1$; if $k=k^{\max}$ go to step 4; else go to step 2
Step 4: Obtain optimal control for $x\in\pazocal{X}$ and $n=0,\ldots,N-1$: $u^{*}_{n}(x)\in\underset{a\in\pazocal{A}_{n}(x)}{\operatorname*{arg\,min}\,}\overline{Q}_{n}(x,a,\theta_{n}^{k^{\max}})$
Calculating the minimum over all actions might be hard to accomplish if the action space is continuous, as in our case. In our numerical experiments, the action space $\pazocal{A}_{n}(x)$ is discretized as for BDP, which has been explained in Section 4.
@@ 6 Numerical Results
In this section, the numerical results obtained with Algorithms 1 and 2, i.e. BDP and Q-learning, are presented and compared with each other. More precisely, we compare the results in terms of accuracy and computational effort of the computed solutions (value function and trajectory). A time horizon of $t_{\text{E}}=120$ hours is selected for the numerical simulation to find the value function and optimal operation for the industrial P2H system during a working week (5 days). The system and algorithm parameters can be found in Tables 1(a) and 1(b), respectively. To keep the results of the proposed algorithms comparable, the same action space discretization is used for both. The evaluation of the minimization over the discretized action space is done by calculating all action values and selecting the action with the minimal value. Penalty costs are applied at terminal time if the TES temperature is below a certain threshold value. In the experiments, the storage must be at least half-full, i.e. the critical value is set to $r_{\text{crit}}=(r_{\text{max}}-r_{\text{min}})/2$. Furthermore, falling below $r_{\text{crit}}$ is penalized with a penalization price $s_{\text{Pen}}=90$/MWh. We do not reward the liquidation of the TES energy and therefore set $s_{\text{Liq}}=0$/MWh. Selling excess energy into the grid is also not allowed, i.e. we set $\delta=0$. For the calculation of the value function, both methods made use of a time-varying state space discretization, see Appendix I. For visualization and convenience, we display the value function for fixed grid points selected from the set $[r_{\text{min}},r_{\text{max}}]\times[2,23.5]\times[15,55]$. This set is chosen such that it is a subset of $\pazocal{X}_{n}$ for all $n=1,\ldots,120$. In order to visualize the results with respect to the three-dimensional state space we will fix some state variables to specific values and plot against the others. In the following, these fixed values are chosen as the corresponding centers of the $r,w$ and $s$-axis.
@@ 6.1 Backward Dynamic Programming
Let us first discuss the results of the value function computed by BDP. Figure 5(a) shows the value function for the initial time step $n=0$ (left) and the terminal costs $n=120$ (right) as a function of TES temperature and wind speed. Obviously, low wind speed and low TES temperature lead to higher expected costs. While the dependence on the storage temperature is almost linear, the wind speed has a significant nonlinear influence on the value function, whereby the latter effect are induced by the WT power curve model. In addition, the dependence of the value function on storage temperature and electricity price as well as on wind speed and electricity price for the initial time is depicted in Figure 5(b) on the left and right, respectively. The electricity price affects the expected costs mostly linearly with respect to the TES temperature, with higher prices leading to higher values. Again, higher storage temperatures reduce costs and compensate for expensive grid electricity. The relation between wind speed and electricity price is also almost linear if we consider changes with respect to the prices. However, in terms of wind speed, the values reflect the nonlinear power curve model.
(a) Value function at initial time $n=0$ (left) and at terminal time $n=120$ (right) in terms of storage temperature and wind speed.
(b) Value function at initial time $n=0$ depending on the storage temperature and electricity price (left) as well as on wind speed and electricity price (right).
Figure 5: BDP: Visualization of the value function for different combinations of state variables.
Second, we analyze a trajectory, starting from the initial state $(R_{0},W_{0},S_{0})=(244.4,4,37)$ and recalling that $R_{0}=r_{crit}$, for the control obtained by BDP as shown in Figure 6(a). As expected, the control aims to charge the TES during periods of high wind power production and/or low prices. Charging only with wind energy can be identified when the HTHP’s electricity consumption (dashed black line) is covered by the available wind energy (green area), while additional power from the grid (blue area) is used to charge the TES when the price falls to small values.
@@
(a) BDP: Visualization of system dynamics and control for one trajectory.
(b) Q-learning: Visualization of system dynamics and control for one trajectory.
@@
For further details on the optimal control $A$, see Figure 7(a), where the optimal decision rule is plotted as a function of time and TES temperature. At each time step $t_{n}=n\Delta t$, the decision rule is calculated for the corresponding TES temperature level (left y-axis) and the values of the seasonalities $\mu_{W}(t_{n})$ and $\mu_{S}(t_{n})$ at that time step. The seasonality functions $\mu_{W}$ (dashed green line) and $\mu_{S}$ (dotted red line) with their corresponding scales are also presented (right y-axis). The colored red and blue areas correspond to the charging and discharging mode of the system with the respective heat flow rate. The white areas represent the idle mode of the system, i.e. no charging or discharging is operated.
It can be seen that the control obtained by BDP captures the functional structure of these seasonalities. This means that when prices are globally low ($\mu_{S}$ takes a global minimum), for example at the hours 25, 50 or 75, charging is the preferred action. Conversely, discharging is performed when prices are high and wind speeds are low. For example, at hours 15, 35 or 60, we can observe that when prices and wind speeds are locally high, the optimal action is to wait and operate in idle mode. Another situation where waiting is optimal, appears when the storage temperature is higher than the penalty temperature $r_{crit}$ (dashed gray line) and prices are at the global minimum of $\mu_{S}$, see e.g. hour 45 to 55. However, when a maximum price is reached, the P2H system operates in discharge mode to compensate for the high electricity prices. Furthermore, when observing the seasonal patterns, it can be seen that the controls for charging, discharging and idle mode are usually centered around the local and global extrema of $\mu_{S}$. In particular, the electricity price appears to have a greater influence on the control than the wind speed (wind energy), as it has a major impact on the operational costs of the P2H system.
(a) BDP: Visualization of the decision rule $u_{n}$.
(b) Q-learning: Visualization of the decision rule $u_{n}$.
@@
(a) Visualization of the value function in terms of storage temperature and wind speed.
(b) Visualization of the value function in terms of electricity price and wind speed.
6.2 Q-Learning
We now compare the results of Q-learning using replay buffer, as shown in Algorithm 2, and BDP based on the computed value function and the corresponding control. For the parameterization of $\overline{Q}(x,a,{\theta_{n}})$ a two-layer neural network with 128 neurons for each layer and ReLU activation functions is used. Since the parameterization is defined globally on the state space, no state discretization is required as for BDP. The state distribution of the initial states $X_{0}=x_{0}$ is chosen as a uniform random distribution on the discretized state space $\widetilde{\pazocal{X}}_{0}$. The Figures 8(a) and 8(b) show the value function of both methods depending on TES temperature and wind speed as well as on wind speed and electricity price, respectively, for the hours 0, 45, 85 and 117. More precisely, in each value function plot of Q-learning, the reference solution of BDP is visualized as gray shaded graph. In addition, the red lines represent a cross-section of the value function, meaning that the value function is fixed in two variables and is visualized in order to provide a more detailed view at the error between both solutions. For a better comparison, the cross-sections (from Figure 8) are also depicted in Figure 9. Compared to BDP, Q-learning is able to capture the same shape of the value function. Especially, for $n$ close to the terminal time, both function approximations differ only slightly. The further we move back in time from the terminal time, the more differences become visible.
Figure 6(b) also confirms that the approximation of the value function by Q-learning is similar to that of BDP. The difference in the control is mainly reflected in the charging and discharging intensity of the TES and thus affects the HTHP’s electricity consumption. Apart from this, the controller aims to charge the TES during times of high wind energy availability or/and low prices and to discharge vice versa. Overall, the comparison shows that the Q-learning control is qualitatively similar to that of BDP. An exception, however, is the control in Figure 6(b), which does not contain any waiting periods. Instead, the charging or discharging periods are generally longer than with the BDP control, as can be seen for in hours 15 to 25. Furthermore, Figure 7(b) provides a more detailed look at the optimal decision rule obtained with Q-learning. Obviously, the charging mode is performed in regions where price seasonality has a global minimum. However, almost every time a peak occurs, the high prices are compensated by discharging the TES. Here too, the role of wind speed appears to be less important than the influence of the electricity price on the control itself. Even though seasonality is taken into account, the overall structure and sequence are not captured as well as with the BDP.
@@ 7 Conclusion
This article presents a mathematical model for the cost-optimal operation of an industrial P2H system. Apart from providing a modeling approach for the stochastic processes that takes into account correlation between wind speed and electricity price, we also calibrated the associated parameters with real data. The resulting discrete-time stochastic optimal control problem is formulated as an MDP and solved using the classical dynamic programming approach as well as modern reinforcement learning techniques, namely Q-learning.
@@
By dropping the assumptions of constant mass flow $\dot{m}$ and waste heat temperature $T^{\text{LT,in}}$, we can extend our model and make it more general. In this case, it is necessary to introduce $T^{\text{LT,in}}$ as an additional state and $\dot{m}$ as an additional control variable, which increases the dimension of the control problem. Even though the classical backward recursion of dynamic programming might become intractable for this extended model due to the curse of dimensionality, we are confident that Q-learning still offers a efficient solution. However, as the dimension of the action space grows, it becomes infeasible to calculate the minimal action by discretization. Appropriate gradient descent methods could be used at the cost that this may slow down the algorithm. Reinforcement learning algorithms such as policy gradients or actor-critic methods as in Sutton2018 might be more suitable for dealing with large (continuous) action spaces, as they already include a way to handle with this minimization step.
@@ Appendix A Nomenclature
HTHP High temperature heat pump
@@
LTHX Low temperature heat exchanger
HTHX High temperature heat exchanger
@@
$m_{S}$ Mass of TES
$c_{S},c_{\text{F}}$ Heat capacity of TES/thermal oil
$A$ Heat flow rate
$B$ Brownian motion
$R$ Storage temperature
$S$ Electricity price of grid
$W$ Wind speed
@@
$N$ Total number of time points
@@
Appendix B Surrogate Models
As described in Walden2023 , the part-load behavior of HTHP and SG is represented by polynomial surrogate models, which are created using process simulation software. The HTHP surrogate models are given by the multivariate polynomials $F_{1}$ and $F_{2}$. Here, $F_{1}$ describes the HTHX outlet temperature $T^{\text{HT,out}}$ and $F_{2}$ the consumed electrical power $P^{\text{H}}$ at a given HTHX inlet temperature $T^{\text{HT,in}}=\tau^{H}$, mass flow $\dot{m}$, LTHX air inlet temperature $T^{\text{LT,in}}=\tau^{L}$ and compressor shaft speed $d$ by
$\begin{split}T^{\text{HT,out}}&=F_{1}(\tau^{H},\dot{m},\tau^{L},d)=95.9612+0.93433\tau^{H}-0.327753\dot{m}+0.0146542\tau^{L}-271.354d\\ &+0.00104853(\tau^{H})^{2}+0.0211819\tau^{H}\dot{m}-0.706122\tau^{H}d-0.00388073\dot{m}\tau^{L}+0.0595068\tau^{L}d\\ &-29.4801\dot{m}d+1.04924\dot{m}^{2}+562.428d^{2}-0.000716825(\tau^{H})^{2}d-2.18172\dot{m}d^{2}-151.476d^{3}\\ &+0.0229386\tau^{H}\dot{m}d+0.881391\dot{m}^{2}d+0.203578\tau^{H}d^{2}-0.0405702\dot{m}^{3}-0.00148575\tau^{H}\dot{m}^{2},\\[4.30554pt] P^{\text{H}}&=F_{2}(\tau^{H},\dot{m},\tau^{L},d)=127.87+2.06342\tau^{H}+2.55723\dot{m}+0.756419\tau^{L}-1164.84d\\ &-1.3829\tau^{L}d-0.0168942\tau^{H}\dot{m}-2.60579\tau^{H}d-0.540713\dot{m}^{2}+13.3204\dot{m}d+1556.66d^{2},\end{split}$ (82)
The surrogate models for SG inlet and outlet temperatures as functions of mass flow read as follows:
Appendix C Model Parameters
The parameters of the model and the methods used to solve the MDP in our numerical experiments, i.e. BDP and Q-learning, are listed in Tables 1(a) and 1(b).
Table 1: Parameters used in the description and application of the physical industrial P2H system, the stochastic processes and the numerical methods.
(a) Parameters used in the overall physical industrial P2H system and the stochastic processes.
Parameters of the Physical P2H System Value Unit

### v3

@@ 5.3 Experience Replay
Q-learning faces the problem that creating trajectories and samples can be a time-consuming task, especially if the dynamics of the system are difficult to simulate or the time horizon is large. This makes it intractable to use generated samples only once and then throw them away when the parameter update is completed. We will use a technique called experience replay LinLong-Ji ; Mnih2013 that solves the problem of wasting generated samples. This uses a so-called replay memory or replay buffer $\pazocal{R}$ with size $\pazocal{N}_{R}\in\mathbb{N}$ to store samples $(x_{n},a_{n},x_{n+1})\in\pazocal{R}$ for $n=0,\ldots,N-1$ and replay (reuse) them as needed in batches $(x_{n}^{j},a_{n}^{j},x_{n+1}^{j}),j=1,\ldots,M$ to update the parameters. Sampling past experience, for example, from a uniform distribution on $\pazocal{R}$, also helps to overcome the exploration-exploitation dilemma. Here, the samples obtained in the early phase of the algorithm are used repeatedly, making the choice of the exploration rate less important for the performance of the algorithm. We summarize the Q-learning method with replay buffer in Algorithm 2.
@@
1: Initialize $(\theta_{n}^{0})_{n=0}^{N-1}$; set the maximum number of iterations $k^{\max}>0$ and $k=0$,
batch size $M$; choose a selection policy $u^{S}=(u_{n}^{S})_{n=0}^{N-1}$
2: Set $n=0$; choose the initial states $x_{0}\in\pazocal{X}$
@@
  end for
  Choose $\alpha_{n}^{k}\in[0,1]$ and update parameters
$\theta_{n}^{k+1}=\theta^{k}_{n}-\alpha_{n}^{k}\frac{1}{M}\sum_{j=1}^{M}\delta_{n}^{j}\nabla_{\theta_{n}}\overline{Q}_{n}(x_{n}^{j},a_{n}^{j},\theta_{n}^{k})$
@@
3: Set $k=k+1$; if $k=k^{\max}$ go to step 4; else go to step 2
4: Obtain optimal control for $x\in\pazocal{X}$ and $n=0,\ldots,N-1$: $u^{*}_{n}(x)\in\underset{a\in\pazocal{A}_{n}(x)}{\operatorname*{arg\,min}\,}\overline{Q}_{n}(x,a,\theta_{n}^{k^{\max}})$
Calculating the minimum over all actions might be hard to accomplish if the action space is continuous, as in our case. In our numerical experiments, the action space $\pazocal{A}_{n}(x)$ is discretized as for BDP, which was explained in Section 4.
@@ 6 Numerical Results
In this section, the numerical results obtained by Algorithms 1 and 2, i.e., BDP and Q-learning, are presented and compared with each other. More precisely, we compare the results in terms of accuracy and computational effort of the computed solutions (value function and trajectory). A time horizon of $t_{\text{E}}=120$ hours is selected for the numerical simulation to find the value function and optimal operation for the industrial P2H system during a working week (5 days). The system and algorithm parameters can be found in Tables 9(c) and 9(d), respectively. To keep the results of the proposed algorithms comparable, the same action space discretization is used for both. The evaluation of the minimization over the discretized action space is done by calculating all action values and selecting the action with the minimal value. Penalty costs are applied at terminal time if the TES temperature is below a certain threshold value. In the experiments, the storage must be at least half-full, i.e., the critical value is set to $r_{\text{crit}}=(r_{\text{max}}-r_{\text{min}})/2$. Furthermore, falling below $r_{\text{crit}}$ is penalized with a penalization price $s_{\text{Pen}}=90$ €/MWh. We do not reward the liquidation of the TES energy and therefore set $s_{\text{Liq}}=0$ €/MWh. Selling excess energy into the grid is also not allowed and we set $\zeta=0$. For the calculation of the value function, both methods made use of a time-varying state space discretization, see Section Appendix I. For visualization and convenience, we display the value function for fixed grid points selected from the set $r_{\text{min}},r_{\text{max}}]\times[2,23.5]\times[15,55]$. This set is chosen such that it is a subset of $\pazocal{X}_{n}$ for all $n=1,\ldots,120$. In order to visualize the results with respect to the three-dimensional state space we will fix some state variables to specific values and plot against the remaining variables. In the following, these fixed values are chosen as the corresponding centers of the $r,w$ and $s$-axis.
@@ 6.1 Backward Dynamic Programming
Let us first discuss the results of the value function computed by BDP, which are presented in Figure 5(a). The graphics on the left and right show the value function for the initial time step $n=0$ and the terminal costs $n=120$, respectively, as a function of TES temperature and wind speed. Obviously, low wind speed and low TES temperature lead to higher expected costs. While the dependence on the storage temperature is almost linear, the wind speed has a significant nonlinear influence on the value function, whereby the latter effect are induced by the WT power curve model. In addition, the dependence of the value function on storage temperature and electricity price as well as on wind speed and electricity price for the initial time is depicted in Figure 5(b) on the left and right, respectively. The electricity price affects the expected costs mostly linearly with respect to the TES temperature, with higher prices leading to higher values. Again, higher storage temperatures reduce costs and compensate for expensive grid electricity. The relation between wind speed and electricity price is also almost linear if we consider changes with respect to the prices. However, in terms of wind speed, the values again reflect the nonlinear power curve model.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: BDP: (a) Value function at initial time $n=0$ (left) and at terminal time $n=120$ (right) in terms of storage temperature and wind speed. (b) Value function at initial time $n=0$ depending on the storage temperature and electricity price (left) as well as on wind speed and electricity price (right).
Second, we analyze a trajectory, starting from the initial state $(R_{0},W_{0},S_{0})=(244.4,4,37)$ with $R_{0}=r_{\text{crit}}$, for the control obtained by BDP as shown in Figure 6(a). As expected, the control aims to charge the TES during periods of high wind power production and/or low prices. Charging that solely uses wind energy can be identified when the HTHP’s electricity consumption (dashed black line) is covered by the available wind energy (green area), while additional power from the grid (blue area) is used to charge the TES when the price falls to small values.
@@
For further details on the optimal control $A$, see Figure 9(a), where the optimal decision rule is plotted as a function of time and TES temperature. At each time step $t_{n}=n\Delta t$, the decision rule is calculated for the corresponding TES temperature level (left y-axis) and the values of the seasonalities $\mu_{W}(t_{n})$ and $\mu_{S}(t_{n})$ at that time step. The seasonality functions $\mu_{W}$ (dashed green line) and $\mu_{S}$ (dotted red line) with their corresponding scales are also presented (right y-axis). The colored red and blue areas correspond to the charging and discharging mode of the system with the respective heat flow rate. The white areas represent the idle mode of the system, i.e., no charging or discharging is operated.
It can be seen that the control obtained by BDP captures the functional structure of the seasonalities. This means that when prices are globally low ($\mu_{S}$ takes a global minimum), for example at the hours 25, 50 or 75, charging is the preferred action. Conversely, discharging is performed when prices are high and wind speeds are low, see hours 20 or 45. For At hours 15, 35 or 60, we can observe that when prices and wind speeds are locally high, the optimal action is to wait and operate in idle mode. Another situation where waiting is optimal, appears when the storage temperature is higher than the penalty temperature $r_{\text{crit}}$ (dashed gray line) and prices are at the global minimum of $\mu_{S}$, see e.g. hour 45 to 55. However, when a maximum price is reached, the P2H system operates in discharge mode to compensate for the high electricity prices. Furthermore, when observing the seasonal patterns, it can be seen that the controls for charging, discharging and idle mode are usually centered around the local and global extrema of $\mu_{S}$. In particular, the electricity price appears to have a greater influence on the control than the wind speed (wind energy), as it has a major impact on the operational costs of the P2H system.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
6.2 Q-Learning
We now compare the results of Q-learning using replay buffer, summarized in Algorithm 2, and BDP based on the computed value function and the corresponding control. Thus, BDP serves as a benchmark, as we can expect accurate results due to the high computational effort. For the parameterization of $\overline{Q}(x,a,{\theta_{n}})$ a two-layer neural network is used with 128 neurons for each layer and ReLU activation functions. Since the parameterization is defined globally on the state space, no state discretization is required as for BDP. The state distribution of the initial states $X_{0}=x_{0}$ is chosen as a uniform random distribution on the discretized state space $\widetilde{\pazocal{X}}_{0}$. The Figures 7(a) and 7(b) show the value function of both methods depending on TES temperature and wind speed as well as on wind speed and electricity price, respectively, for the hours 0, 45, 85 and 117. More precisely, in each value function plot of Q-learning, the reference solution of BDP is visualized as gray shaded graph. In addition, the red lines represent a cross-section of the value function, meaning that the value function is fixed in two variables and is visualized in order to provide a more detailed view at the error between both solutions. For a better comparison, the cross-sections (from Figure 7) are also depicted in Figure 8. Compared to BDP, Q-learning is able to capture the same shape of the value function. Especially, for $n$ near the terminal time, both function approximations differ only slightly. The further we move back in time from the terminal time, the more differences become visible.
Figure 6(b) also confirms that the approximation of the value function by Q-learning is similar to that of BDP. The difference in the control is mainly reflected in the charging and discharging intensity of the TES and thus affects the HTHP’s electricity consumption. Apart from this, the controller aims to charge the TES during times of high wind energy availability or/and low prices and to discharge vice versa. Overall, the comparison shows that the Q-learning control is qualitatively similar to that of BDP. However, an exception is the control in Figure 6(b), which almost does not contain any waiting periods. Instead, the charging or discharging periods are generally longer compared to the BDP control, which can be seen in hours 15 to 25. Furthermore, Figure 9(b) provides a more detailed look at the optimal decision rule obtained with Q-learning. Obviously, the charging mode is performed in regions where price seasonality has a global minimum. However, almost every time a peak occurs, the high prices are compensated by discharging the TES. Again, the role of wind speed appears to be less important than the influence of the electricity price on the control itself. Even though seasonality is taken into account, the overall structure and sequence are not captured as well as with the BDP.
@@ 7 Conclusion
This work presents a mathematical model for the cost-optimal operation of an industrial P2H system. Apart from providing a modeling approach for the stochastic processes that takes into account correlation between wind speed and electricity price, we also calibrated the associated parameters with real-world data (see Supplementary Material G). The resulting discrete-time stochastic optimal control problem is formulated as an MDP and solved using the classical dynamic programming approach as well as modern reinforcement learning techniques, namely Q-learning.
@@
By dropping the assumptions of constant mass flow $\dot{m}$ and waste heat temperature $T^{\text{LT,in}}$, we can extend our model and make it more general. In this case, it is necessary to introduce $T^{\text{LT,in}}$ as an additional state and $\dot{m}$ as an additional control variable, which increases the dimension of the control problem. Even though the classical backward recursion of dynamic programming might become intractable for this extended model due to the curse of dimensionality, we are confident that Q-learning still offers an efficient solution. However, as the dimension of the action space grows, it becomes infeasible to calculate the minimal action by discretization. Appropriate gradient descent methods could be used, at the cost that this may slow down the algorithm. Reinforcement learning algorithms such as policy gradients or actor-critic methods as in Sutton2018 might be more suitable for dealing with large (continuous) action spaces, as they already include a way to handle this minimization step.
@@ Appendix A Nomenclature
HTHP High-temperature heat pump
@@
LTHX Low-temperature heat exchanger
HTHX High-temperature heat exchanger
@@
$m_{\text{s}}$ Mass of TES
$c_{\text{p,s}},c_{\text{p,f}}$ Heat capacity of TES/thermal oil
$d$ Rotational speed
@@
AA Heat flow rate
$B$ Brownian motion
$X$ State process
$Y$ Ornstein-Uhlenbeck process
$R$ Storage temperature
$S$ Electricity grid price
$W$ Wind speed
$Z$ Random disturbance
$\widehat{Z}$ Quantizer for $Z$
@@
NN Total number of time points
$n_{\text{H}}$ Number of HTHPs running in parallel
@@
$\pazocal{T}$ Transition operator
$a$ Control/action variable
$C_{n}$ Running cost
$G$ Terminal cost
$J_{n}$ Performance criterion
$V_{n}$ Value function
$Q_{n}$ State action function
$\theta$ Parameter vector
@@
Appendix B Model Parameters
The parameters of the model and the methods used to solve the MDP in our numerical experiments, i.e., BDP and Q-learning, are listed in Tables 9(c) and 9(d).
Table 1: Parameters used in the description and application of the industrial P2H system, the stochastic processes and the numerical methods.
Parameters of the P2H System Value Unit